OpenStudy (anonymous):

what is the solution to the system? use substitution. y = x + 6 y = 2x a. 6,12 b. -12,-6 c. -6,-12 d. 2,4

OpenStudy (jojo4eva):

the answer would be A. 6,12

OpenStudy (anonymous):

-_-

OpenStudy (anonymous):

x=6 y=12

OpenStudy (jojo4eva):

12=(6)+6 12=12 12=2(6) 12=12 So A. 6,12 is your final answer